Mary is in contract negotiations with a publishing house for her new novel. She has two options. She may be paid $100000 up front, and receive royalties that are expected to total $36000 at the end of each of the next 5 years. Alternatively, she can receive $200000 up front and no royalties. Which of the following investment rules would indicate that she should take the former deal, given a discount rate of 10%? Rule I: The Net Present Value rule; Rule II: The Payback Rule with a payback period of 2 years;"

A) Rule I only

B) Rule II only

C) Rule I and II

D) None
Answer #1

NPV = Present value of cash inflows – Present value of cash outflows

= Option 1 = 100,000 + 36000*PVAF(10%, 5 years)

= 100,000 + 36,000*3.791

= $236,476

Option 2= 200,000

Hence, Option 1 is better as per NPV since higher NPV

Payback period = Difference in initial amount/Annual inflow

= 100,000/36000

= 2.78 years

Hence, Option 1 is not acceptable as per payback period method since higher than 2 years

Hence, the answer is RULE I only

i.e. A